OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use 3.14 for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

v=(1/3) pi r^2 h

OpenStudy (anonymous):

I got that same exact answer last time ... Can you show me how to do it? or at least Give me the answer?

OpenStudy (anonymous):

300=(1/3)*49 (3.14) h ===> h = 5.849 km

OpenStudy (anonymous):

Are you sure thats the answer?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is that rounded to the nearest tenth??

OpenStudy (anonymous):

5.8

jimthompson5910 (jim_thompson5910):

V=(1/3)pi*r^2*h 3V=pi*r^2*h (3V)/(pi*r^2)=h h=(3V)/(pi*r^2) h=(3(300))/(pi*(7)^2) h=(3(300))/((3.14)*(7)^2) h=(3(300))/((3.14)*(49)) h=(900)/((3.14)*(49)) h=(900)/(153.86) h=5.84947354738074 So the height is approximately 5.8494735 km

jimthompson5910 (jim_thompson5910):

to the nearest tenth, the height is 5.8 km

OpenStudy (anonymous):

Jim Thompson thanks :) Please look at the lastest questions Ive asked